OpenStudy (anonymous):

cost of hat: $10.50 Markup:10% Discount:40% Tax:5% help please with the simple interest formula.

OpenStudy (anonymous):

\[10.50\times (1.10)\times (.60)\times (1.05)\] \(1.10\) gives the 10% markup, \(.60\) takes off 40% and \(1.05\) increases by 5%

OpenStudy (anonymous):

where do you get the decimals from?

OpenStudy (anonymous):

10% =.1 5%=.05 40%=.4

OpenStudy (anonymous):

to increase a number by 10% you are taking 110% of a number, i.e. 1.10 to decrease a number by 40% you are taking 60% of that number, i.e. .6

OpenStudy (anonymous):

so would it be $10.50(1.10)(1.4)(1.05)=16.97 ?

OpenStudy (anonymous):

no because the 40% decrease should mean multiply by \(100\%-40\%=60\%=.6,\) not 1.4

OpenStudy (anonymous):

hence my first post of \[10.50\times (1.10)\times (.60)\times (1.05)\]

OpenStudy (anonymous):

one more question lol. do we always subtract by 100 with the discount? & thank you very much for your help. really appreciate it, it's helping me alot.

OpenStudy (anonymous):

@satellite73